Cd2+(aq) + 4CH3NH2(aq)≤≥ [Cd(CH3NH2)4]
2+(aq) Kstab = 3.6 × 106 equilibrium I

The concentration of Cd2+(aq) is 1.00×10^-4 moldm-3
Calculate the concentration of CH3NH2(aq) needed to reduce the concentration of Cd2+(aq) in this dilute solution by a factor of one thousand.
1
Expert's answer
2020-11-04T14:14:36-0500

x mol/dm3 of CH3NH2(aq) was added

1*10-7 mol/dm3 Cd2+(aq) left

9.99*10-5 mol/dm3 of Cd2+(aq) and 3.996*10-4 mol/dm3 of CH3NH2(aq) reacted

9.99*10-5 mol/dm3 of [Cd(CH3NH2)4]2+(aq) formed

(x-3.996*10-4) mol/dm3 of CH3NH2(aq) left

3*106 = (9.99*10-5)/(1*10-7*(x-3.996*10-4)4)

x-3.996*10-4 = 0.1351

x=0.1355 mol/dm3 - concentration of CH3NH2(aq) needed